# Analyzing Local Representations of Self-supervised Vision Transformers

### 31 Dec 2023 v1

@@ Abstract
In this paper, we present a comparative analysis of various self-supervised Vision Transformers (ViTs), focusing on their local representative power. Inspired by large language models, we examine the abilities of ViTs to perform various computer vision tasks with little to no fine-tuning. We design evaluation framework to analyze the quality of local, i.e. patch-level, representations in the context of few-shot semantic segmentation, instance identification, object retrieval and tracking. We discover that contrastive learning based methods like DINO produce more universal patch representations that can be immediately applied for downstream tasks with no parameter tuning, compared to masked image modeling. The embeddings learned using the latter approach, e.g. in masked autoencoders, have high variance features that harm distance-based algorithms, such as k-NN, and do not contain useful information for most downstream tasks. Furthermore, we demonstrate that removing these high-variance features enhances k-NN by providing an analysis of the benchmarks for this work and for Scale-MAE, a recent extension of masked autoencoders. Finally, we find an object instance retrieval setting where DINOv2, a model pretrained on two orders of magnitude more data, performs worse than its less compute intensive counterpart DINO.
@@ 1 Introduction
Recent advances in Natural Language Processing gave birth to universal models that after large-scale pretraining can perform various language-related tasks without task-specific fine-tuning. Large language models [3, 18, 12] based on self-supervised transformers achieve competitive performance on tasks like translation, question answering, and commonsense reasoning with prompting or by in-context learning with just a few examples.
Self-supervised transformers are also getting increasingly popular in computer vision. Two radically different self-supervised learning paradigms have demonstrated good performance for Vision Transformers (ViTs): those based on contrastive learning (e.g. MOCO [14] or DINO [4]), and those based on masked image modeling (e.g. MAE [15] or SimMIM[29]). The question of whether these models possess universal capabilities, similar to those seen in NLP models, for computer vision tasks remains unanswered.
As ViTs do not have text inputs, it is non-trivial to assess their zero-shot capabilities for downstream tasks. Most ViTs produce one embedding vector for the entire image (usually the [CLS] token) and one embedding for each local patch. In this paper we focus on few-shot capabilities of ViTs for vision tasks that require locality awareness, like image segmentation and object tracking. We propose few-shot evaluation methods that leverage patch representations. To minimize task-specific parameter tuning we use two approaches: k-nearest-neighbors with no tuned parameters and linear probing with a single layer of trainable parameters. The power of global image representations from pretrained ViTs for image-level tasks like image classification are relatively well explored in literature, e.g. [4].
Fine-tuning the entire backbone in addition to relatively large task-specific “heads” still gives superior performance in segmentation and tracking. The analysis of such models is beyond the scope of this paper as their good performance is not just from self-supervised pretraining, but also strongly affected by the head architecture and the data used for fine-tuning. The focus of this work is on the inherent capabilities of self-supervised ViTs that can be exposed by using just a few labeled samples.
We show that while masked image modeling produces backbones with good fine-tuning performance, the frozen, pretrained patch embeddings are far inferior to the ones learned by contrastive methods for nearest neighbor methods. We dive deep into this phenomenon and identify roughly 200 dataset-agnostic features in the embedding space that, counterintuitely, contain no useful information for the downstream tasks we have considered, while having the highest variances among all features. Removing those features improves k-NN performance for most tasks.
@@
  • We design an evaluation framework along with few-shot datasets to analyze inherent power of pretrained Visual Transformers for locality aware tasks. We analyze and compare five representative ViTs using our framework on three tasks: patch classification, instance and fine-grained retrieval, and object association in video frames.
@@
  • We demonstrate that the removal of these features is beneficial not only for our proposed benchmarks but also for benchmarks presented in other studies. For instance, Scale-MAE [25], which assesses a MAE-like network trained on aerial images, evaluates its performance across various resolutions and compares it against other state-of-the-art algorithms using k-NN on computed representations (either averaged for all patches or the CLS token). We observe the same issue of high-variance features, and after removing 200 such features, we showcase a superior performance.
@@ 2 Related work
Park and Kim [20] analyze and demonstrate several properties of multi-head self-attentions (MSA) and ViTs. They show that MSAs flatten the loss landscape to alleviate the issue of its non-convexity. They also observe that MSAs and convolutional layers complement each other, showing that MSAs function as low-pass filters, while convolutional layers act as high-pass filters.  Park et al. [21] analyze the differences between the ViT methods that are based on contrastive learning (CL) and masked image modeling (MIM) and compare their performance of downstream tasks. They demonstrate that CL captures longer-range global patterns, such as object shapes, more effectively than MIM methods. Secondly, they demonstrate that CL-based approaches are more shape-oriented, whereas MIM-based approaches are more texture-oriented.  Raghu et al. [24] provide an analysis of the internal representation structure of ViTs and CNNs on several image classification benchmarks. They demonstrate that ViTs have a more uniform representation across the layers of the network compared to CNNs. These differences are mostly explained by the role of self-attention, which allows early aggregation of information, and ViT residual connections, which also propagate features from lower to higher levels.
Other works have focused on analyzing the robustness of ViTs.  Bhojanapalli et al. [2] investigate the robustness of ViT models to input and model perturbations for image classification.  Bhojanapalli et al. [2] demonstrate that transformers are robust to the removal of almost any single layer and that when pre-trained on a sufficiently large dataset, ViTs demonstrate not worse results than the ResNet counterparts across various perturbations.  Paul and Chen [22] analyze the robustness of ViTs against several common corruptions, perturbations, distribution shifts, and natural adversarial examples. They also analyze and demonstrate the superior robustness of ViTs in various aspects, such as masking, energy/loss landscape analysis, and sensitivity to high-frequency artifacts on robust classification datasets. [17] investigate the robustness of Transformers to severe occlusions, perturbations, and domain shifts in classification tasks. Their findings demonstrate that ViTs exhibit significantly less bias towards local textures compared to CNNs.
@@
While not precisely the same phenomenon, a somewhat similar occurrence of artifacts in feature maps is analyzed in [8]. The study demonstrates that for supervised and self-supervised Vision Transformers (VITs), certain artifact high-norm tokens emerge during inference, primarily in low-informative background areas of images. It is worth noting that our observation for MAE is the presence of features with high variance, whereas [8] observes that certain tokens exhibit artifacts.
@@
3 Can Transformers Recognize Semantics of Patches?
@@
Figure 1: mIoU of patch classifiers on various subsets of Cityscapes dataset. The evaluation is performed on the same validation set. The size of the circles are proportional to the number of images in the training set (36 images, 72 images and 144 images).
To analyze the local representations of the ViT models, we first study their ability to perform patch-wise classification. To this end, we set up a few-shot patch classification experiment on the Cityscapes dataset [6].
@@
Few-shot subset of Cityscapes. Cityscapes [6] training set consists of 2975 pictures taken in 18 cities. Unless otherwise stated, we use a training dataset consisting of 4 images per city (72 images in total). We later also explore the impact of the number of training sampled by increasing or decreasing this number. The original validation set of the Cityscapes dataset has $500$ images, containing images from 3 different cities. For our analyses, we take $10$ images per city, resulting in a total of $30$ images. We convert the pixel-dense segmentation labels of the Cityscapes dataset to patch-level classes by selecting the most common class within each patch. We evaluate the quality of representations by measuring pixel accuracy and segmentation mIoU.
As the ViT models used in our analyses take as an input an image of size $224\times 224$, we tile the images of size $1024\times 2048$ into $256\times 256$ tiles and treat each tile as a separate image. The tiles are resized to $224\times 224$ and passed to the pre-trained transformers. We extract and store the patch representations of the corresponding transformer for all the patches in all the images (training and validation) and for all the ViTs.
k-NN and linear probing. We use two simple classifiers to analyze the underlying representations: k-NN with $k=1$ and fitting a linear softmax classifier. Both are trained on the patch representations of the few-shot training set. The motivation for taking these two basic methods is to understand whether the patches of a given object category cluster together or are linearly separable from other object categories in the representation space. In case of MAE, we follow recommendations from [15] (which on its turn refers to [10]) and apply batch normalization to the extracted features before the linear layer. We note that we get almost 2.5×2.5$\times$2.5 × better results for linear probing with BatchNorm. We do not use BatchNorm for SimMIM, as adding it worsens the performance.
The size of the training set matters for linear models, less for k-NN. In Fig. 1 we illustrate the performance of the k-NN and linear classifiers for different sizes of the training dataset. We first find the linear models achieve generally get better results than the k-NN, especially with more training data. However, for the DINO versions and the supervised ViT, the k-NN and linear classifiers still perform comparable to each other. Instead, the MAE representations yield surprisingly bad k-NN performance. While its linear results are only slightly inferior to its DINO counterpart, the k-NN classifier leads to a remarkable about 4 times worse performance.
Distinct layer-wise behavior across ViTs. Next, we perform a detailed analysis of the k-NN patch classification performance for representation extracted at different layers of the network in Fig. 2. Supervised ViT and DINO perform strikingly similar. The performance slowly improves from the first layer to the 8-th layer and then saturates. There is a slight drop of performance in the last two layers, which is more visible in case of Supervised ViT.
DINOv2’s behavior is quite different. In the first layers its performance is worse than DINO. For the easiest five object categories (road, vegetation, sky, car, building) the performance catches up in the last layer. For the other, harder object categories DINOv2 catches up with DINO and Supervised ViT in the 10th layer, and then significantly outperforms them in the 11th and 12th layers. In particular, IoU of bus category jumps from 0.059 on the 9th layer to 0.729 on the 12th layer. The advantage of DINOv2 thus mostly stem from the harder categories.
Again, we observe poor k-NN performance in case of the MAE. SimMIM, another ViT trained to reconstruct images, performs better than MAE, but only in the middle layers. The difference is more significant for easier object categories. The quality of the last three layers is similar to MAE.
We applied three sets of corruptions to the same small subset of the validation set of Cityscapes. First, we blur out the images by applying four sizes of the blur kernel: $10\times 10$, $20\times 20$, $30\times 30$ and $40\times 40$. Second, we add four different levels of Gaussian noise (with mean 00 and standard deviations $10,20,30,$ and $40$). At last, we add frequency-based random noise (with $40$ standard deviation) to the images the following way. We generate a Gaussian noise of the same shape as the source image, convert it into the frequency domain, keep the frequencies only inside one of the four narrow bands, convert the noisy image back to image pixels, and add it to the original image. Fig. 10 in Appendix visualizes this process. We feed these degraded images to the ViTs and check the robustness of the methods for the different types and levels of degradations. The first row of Fig. 3 demonstrates the results of kNN for blurred, additive Gaussian noise and added frequency-based random noise for each color, respectively.
One can claim that the robustness to various degradations may be attributed to the augmentations used during training. Specifically, models that employ color enhancement augmentations are assumed to exhibit higher resilience against those types of degradations (blur, Gaussian noise, etc). However, according to [15] color jittering-based augmentation degrades its performance. This suggests a potential trade-off between performance and robustness to degradations. Due to computational limitations, further investigation of this hypothesis is left for future work.
Figure 3: Robustness of linear models for segmentation on Cityscapes. For every model (differing by colors) we have five dots, each representing one level of degradation. The y-axis shows the mIoU on corrupted images. The top most dot in each vertical line corresponds to the original images. Grey diagonal lines indicate levels of equal relative drop in mIoU.
4 Enhancing k-NN Accuracy in Masked Image Modeling
@@
In this section, we tackle a challenge with MIM-based methods in k-NN tasks. The problem revolves around features that have significantly higher variance, compared to other features. We suggest a solution for this issue and then take a closer look at what information these features hold.
4.1 Why MIM-based models have poor k-NN performance.
@@
Throughout our experiments we discovered that MAE’s patch embeddings perform reasonably well with linear probing, but fail with k-NN. We hypothesize that the patch embeddings of MAE might have high variance in some dimensions which could drastically increase the distances among same-category patches and therefore harm k-NN, while not affecting the performance of (possibly small-margin) linear models.
Variances of MAE and SimMIM features are extremely diverse. We computed the variance for each of the features for all models. In Fig. 4(a), we plot the variances of all 768 features in a decreasing order. We observe that the variances throughout all features of Supervised ViT (and DINO) are relatively uniform. However, for MAE and SimMIM, there are several features with very high variances and there is a long tail of close to zero variance features.
4.2 Simple Remedy
Improving k-NN performance for MAE. We removed $m$ features with the highest variances and measured few-shot segmentation performance using k-NN and linear probing with the shortened embeddings. As seen in Fig. 4(b), with just $m=10$ features removed, MAE’s k-NN performance jumped from $0.058$ to $0.170$, without sacrificing the performance of linear probing. The performance of k-NN kept increasing up to $0.295$ at $m=200$. Afterwards, the scores for both k-NN and linear probing started to decline. This finding implies that around a quarter of features of MAE’s embeddings are not useful for patch-level image segmentation for neither linear models nor k-NNs. On the other hand, these features comprise nearly all of the variance of the embeddings.
Improving k-NN performance for Scale-MAE [25]. To validate our hypothesis, we selected a recent MIM-based algorithm that employs k-NN for performance comparison against state-of-the-art methods. We conducted experiments by removing its $m=200$ high-variance features.
Scale-MAE is trained in a way that it has a meaningful [CLS] token, which can be used as a representation of an image along with the mean of patch tokens. The authors used [CLS] token representations for image classification using k-NN and showed that it outperforms MAE. We reproduced the results and confirmed that the mean of patch vectors is indeed worse than [CLS]. However, after removing the high-variance features, mean patch representations outperform [CLS] across almost all values of ground sampling distance (GSD) in the UCMerced and RESISC datasets. The results are presented in Fig. 4. We observe that removing 200 high-variance features from [CLS] does not change the results significantly; however, once we use the mean vectors across the image patches, and after removing 200 high-variance features from the patch representations, they outperform the proposed methods that utilize [CLS] representation.
Figure 4: Comparison between the original Scale-MAE and our proposed variants, where we remove 200 high-variance features, using k-NN on three datasets for various ground sampling distance (GSD) values.
4.3 What information do these features hold?
One hypothesis is that these features are necessary for identifying specific instances of objects among the same category, or for distinguishing fine-grained categories of objects. In Section 5 we provide negative evidence for this hypothesis: removing high-variance features improves retrieval performance so much that the negative effect of losing some instance-specific information, if true, is not detectable.
High variance features do not store global context. Another hypothesis is that the high variance features might contain global information about the image not directly relevant to the semantic class of the current patch. An example of a global information is the number of pixels of each class in the entire image. To test whether patch representations contain such information, we use L2-regularized linear regression to predict those numbers (normalized by the total number of pixels) for the three most common classes (road, building, vegetation) from the representation of a central patch. We report $R^{2}$ for each class and average them. We see that MAE embeddings without top 200 high-variance features have the same predictive power (0.672) as the full MAE embeddings (0.676). It implies that the high variance features do not contain additional information about the global context not present in the other features. Furthermore, the features with the highest variances are less predictive (0.550) than a random subset of 200 features (0.666); see Table 6 of Appendix.
Feature normalization has a similar effect for MAE, but not for SimMIM. We created another version of MAE embeddings by applying the pretrained BatchNorm layer from the linear model, before passing them to k-NN classifier. It helped to improve k-NN performance by almost the same amount as removing $m=200$ features. It improved SimMIM’s k-NN performance as well, but linear probing worsened significantly. We conclude that feature normalization is an alternative but not an identical strategy to feature removal for minimizing the negative impact of high-variance features.
High variance features of MAE are relatively stable across datasets. We identify top 200 high-variance features of patch representations extracted from Cityscapes, ADE20K and FAIR1M datasets. 196 of these features are shared among Cityscapes and ADE20K, while 192 of them are shared among Cityscapes and FAIR1M.
@@ 4.3 What information do these features hold?
(a) Variances of each of the 768 dimensions of embeddings of various ViTs sorted in decreasing order. Note that both axis are logarithmic.
@@
5 At Which Granularity Transformers Detect Objects?
In Sec. 3 we analyzed the performance of ViTs for the object-level few-shot semantic segmentation task. This involves assigning a class to each patch that contains predefined objects, such as airplanes and cars. This led to the next question of whether ViTs can learn to distinguish between objects of the same category but of different types (such as types of cars or types of airplanes) or whether ViTs can perform well in distinguishing the same exact object when other instances of it are available.
Methodology. We take FAIR1M [27], a large dataset of satellite imagery designed for fine-grained object detection. Note that none of the ViTs we tested were pretrained on satellite imagery (even DINOv2, as far as we can tell). The objects in FAIR1M are annotated with respect to 5 supercategories: airplane, ship, vehicle, court, and road and 37 fine-grained categories (types of airplanes, types of ships, etc.). The annotations are in the form of rotated bounding boxes (without pixel precision). However, to the best of our knowledge, all object instances appear on only one image. For this set of experiments, we include a recent extension of MAE called Scale-MAE [25], specifically trained for satellite imagery. To test our hypothesis regarding high-variance features affecting distance-based metrics for MAE, we also use MAE-200 and Scale-MAE-200. These variants remove the 200 high-variance features from the representation vectors.
The images have varying sizes, usually larger than $1000\times 1000$px. We cropped all of them to $224\times 224$px tiles and kept only 8 images per each fine-grained category. This guarantees that our dataset contains at least 8 instances of each category, but in fact it contains many more instances of the common objects. Table 3 in Appendix lists dataset statistics. We have 295 images with 196 patches each for ViT-B/16 models, and 256 patches each for DINOv2.
Next, we created transformed versions of all images. For the first set of experiments, we applied diagonal shifts by $1,2,3,$ and $4$ pixels. For the second set of experiments we rotated the images by $5,$ $10,$ $15,$ and $20$ degrees counter clockwise. For the third set of experiments, we scaled the images by factors $0.8$, $0.9$, $1.1$ and $1.2$. For the fourth and fifth sets of experiments, we applied blur and Gaussian noise degradations. We computed the patch representations of all these images. This creates many images with the same instances of objects.
For the experiments involving image shifting, rotation, and scaling, the first level of evaluation, called ’same-patch,’ is not straightforward because it is not immediately clear which patch corresponds to the original one. We define the corresponding patch as the one containing the center point of the rotated patch. We remark, that several patches near the corners of the rotated tile may lack a corresponding patch, placing an upper bound on same-patch retrieval accuracy. The results, along with the upper bounds, are shown in Fig. 6. We observe, that for rotation and scaling all models exhibit significantly less robustness even with a 5-degree rotation and a small amount of scaling compared to the highest blur radius or noise level attempted. The order of model performance remains consistent with other image transformations: DINO performs the best, followed by MAE (as well as Scale-MAE) with 200 high-variance features removed, followed by DINOv2 and Supervised ViT, with MAE as a distant outlier. It is worth noting that Scale-MAE outperforms MAE significantly; however, removing its 200 high-variance features also notably improves its performance.
DINO is the most robust one. Among the ViTs we tested, DINO is the most robust one across all transformations. Surprisingly, DINOv2 is less robust by all metrics, and performs similar to Supervised ViT. To verify whether this disadvantage of DINOv2 comes from the patch-level loss term or the scale of the model and the dataset, we perform the same analysis with iBOT embeddings. iBOT performs even slightly better than DINO (Fig. 11 in Appendix), which means that patch-level loss cannot be blamed.
Retrieving patches from other images. We notice that most of the retrieved patches come from the same image tile. One possible explanation is that the patch embeddings contain image-level information. We repeat this experiment with the patches of the original image removed from the set of available patches. In this setting the closest patch can be of the same fine-grained category, same supercategory, incorrect supercategory and can be a background patch. Here DINOv2 takes the lead, Supervised ViT and DINO perform slightly worse, and MAE performs poorly (Fig. 7).
MAE’s performance can be improved. We repeated this experiment with the shortened embeddings of MAE and Scale-MAE (200 highest-variance features removed). This version of MAE and Scale-MAE performs significantly better than the original ones (MAE and Scale-MAE) and outperforms both DINOv2 and Supervised ViT, and falls short only of DINO. This result implies that removing the high-variance features of MAE embeddings not only helps in semantic segmentation, but also in the identification of specific object instances in transformed images. On the other hand, this adds to the evidence that the high-variance features do not hold any unique information necessary for identifying the instances, as seen by almost ideal performance on the patches with small transformations (Fig. 6).

### 21 Mar 2024 v2

@@ Abstract
In this paper, we present a comparative analysis of various self-supervised Vision Transformers (ViTs), focusing on their local representative power. Inspired by large language models, we examine the abilities of ViTs to perform various computer vision tasks with little to no fine-tuning. We design evaluation framework to analyze the quality of local, i.e. patch-level, representations in the context of few-shot semantic segmentation, instance identification, object retrieval and tracking. We discover that contrastive learning based methods like DINO produce more universal patch representations that can be immediately applied for downstream tasks with no parameter tuning, compared to masked image modeling. The embeddings learned using the latter approach, e.g. in masked autoencoders, have high variance features that harm distance-based algorithms, such as k-NN, and do not contain useful information for most downstream tasks. Furthermore, we demonstrate that removing these high-variance features enhances k-NN for MAE, as well as for its recent extension Scale-MAE. Finally, we find an object instance retrieval setting where DINOv2, a model pretrained on two orders of magnitude more data, falls short of its less compute intensive counterpart DINO.
@@ 1 Introduction
Recent advances in Natural Language Processing gave birth to universal models that after large-scale pretraining can perform various language-related tasks without task-specific fine-tuning. Large language models [3, 20, 12] based on self-supervised transformers achieve competitive performance on tasks like translation, question answering, and commonsense reasoning with prompting or by in-context learning with just a few examples. Self-supervised transformers are also getting increasingly popular in computer vision. Two radically different self-supervised learning paradigms have demonstrated good performance for Vision Transformers (ViTs): those based on contrastive learning (e.g. MOCO [14] or DINO [4]), and those based on masked image modeling (e.g. MAE [15] or SimMIM [31]). The question of whether these models possess universal capabilities, similar to those seen in NLP models, for computer vision remains unanswered.
As ViTs do not have text inputs, it is non-trivial to assess their zero-shot capabilities for downstream tasks. Most ViTs produce one embedding vector for the entire image (usually the [CLS] token) and one embedding for each local patch. In this paper we focus on few-shot capabilities of ViTs for vision tasks that require locality awareness, like image segmentation and object tracking. We propose few-shot evaluation methods that leverage patch representations. To minimize task-specific parameter tuning we use two approaches: k-NN and linear probing with a single layer of trainable parameters. The power of global image representations from pretrained ViTs for image-level tasks like image classification are relatively well explored in literature, e.g. [4].
Fine-tuning the entire backbone in addition to relatively large task-specific “heads” still gives superior performance in segmentation and tracking. The analysis of such models is beyond the scope of this paper as their good performance is not just from self-supervised pretraining, but also strongly affected by the head architecture and the data used for fine-tuning. The focus of this work is on the inherent capabilities of self-supervised ViTs that can be exposed by using just a few labeled samples. We show that while masked image modeling produces backbones with good fine-tuning performance, the frozen, pretrained patch embeddings are far inferior to the ones learned by contrastive methods for nearest neighbor methods. We dive deep into this phenomenon and identify roughly 200 dataset-agnostic features in the embedding space that, counterintuitely, contain no useful information for the downstream tasks we have considered, while having the highest variances among all features. Removing those features improves k-NN performance for most tasks.
@@
  • We design an evaluation framework along with few-shot datasets to analyze inherent power of pretrained ViTs for locality aware tasks. We analyze and compare five ViTs using our framework on three tasks: patch classification, instance/fine-grained retrieval, and object association in video frames.
@@
  • We demonstrate that the removal of these features is beneficial not only for our proposed benchmarks but also for benchmarks presented in other studies. For instance, Scale-MAE [27], which assesses a MAE-like network trained on aerial images, evaluates its performance across various resolutions and compares it against other state-of-the-art algorithms using k-NN on computed representations. We observe the same issue of high-variance features, and after removing 200 such features, we showcase a superior performance.
@@ 2 Related work
[22] analyze and demonstrate several properties of multi-head self-attentions (MSA) and ViTs. They show that MSAs flatten the loss landscape to alleviate the issue of its non-convexity. They also observe that MSAs and convolutional layers complement each other, showing that MSAs function as low-pass filters, while convolutional layers act as high-pass filters.  [23] analyze the differences between the ViT methods that are based on contrastive learning (CL) and masked image modeling (MIM) and compare their performance of downstream tasks. They demonstrate that CL captures longer-range global patterns, such as object shapes, more effectively than MIM methods. Secondly, they demonstrate that CL-based approaches are more shape-oriented, whereas MIM-based approaches are more texture-oriented. [26] provide an analysis of the internal representation structure of ViTs and CNNs on several image classification benchmarks. They demonstrate that ViTs have a more uniform representation across the layers of the network compared to CNNs. These differences are mostly explained by the role of self-attention, which allows early aggregation of information, and ViT residual connections, which also propagate features from lower to higher levels.
Other works have focused on analyzing the robustness of ViTs.  [2] investigate the robustness of ViT models to input and model perturbations for image classification.  [2] demonstrate that transformers are robust to the removal of almost any single layer and that when pre-trained on a sufficiently large dataset, ViTs demonstrate not worse results than the ResNet counterparts across various perturbations.  [24] analyze the robustness of ViTs against several common corruptions, perturbations, distribution shifts, and natural adversarial examples. They also analyze and demonstrate the superior robustness of ViTs in various aspects, such as masking, energy/loss landscape analysis, and sensitivity to high-frequency artifacts on robust classification datasets. [19] investigate the robustness of Transformers to severe occlusions, perturbations, and domain shifts in classification tasks. Their findings demonstrate that ViTs exhibit significantly less bias towards local textures compared to CNNs.
@@
While not precisely the same phenomenon, a similar occurrence of artifacts in feature maps is analyzed in [8]. The study demonstrates that for supervised and self-supervised Vision Transformers (VITs), certain artifact high-norm tokens emerge during inference, primarily in low-informative background areas of images. It is worth noting that our observation for MAE is the presence of features with high variance, whereas [8] observes that certain tokens exhibit artifacts.
Figure 1: mIoU of patch classifiers on various subsets of Cityscapes dataset. The evaluation is performed on the same validation set. The size of the circles are proportional to the number of images in the training set (36, 72, and 144).
@@
3 Evaluation Frameworks
@@
To fairly assess the effectiveness of existing ViTs, and compare them with our proposed MAE-200 with improved representations Sec. 4, it is important to establish a comprehensive evaluation framework. The objective of our proposed evaluation framework is assessing the effectiveness of existing SSL methods for several downstream tasks, using linear probing and k-NN. We consider 3 different setups for evaluations. The first setup includes analyzing the local representations of the ViTs, by performing patch-wise classification. To this end, we set up a few-shot patch classification experiment on the Cityscapes dataset [6]; see Sec. 5. The second setup analyzes the performance of ViTs for object-level few-shot semantic segmentation task. This task involves classifying patches with predefined objects like airplanes and cars. We explore whether ViTs can distinguish between different types within the same category or effectively differentiate between instances of the same object with multiple occurrences. We use the FAIR1M dataset [29]; see Sec. 5. In the third setup, we assess the ability of ViTs to track objects. Object tracking involves identifying the same object instance across frames in a video. We analyze the robustness of patch embeddings over time, especially as objects undergo appearance changes. We utilize the track-validation set of MOT 2020 images, a subset of the BDD-100K dataset [32]; see Sec. 5. At last, in Sec. 6, we discuss additional numerical experiments under different setups to support our claims.
@@
Cityscapes Dataset [6]. The training set contains 2975 pictures taken in 18 cities. Unless otherwise stated, we use a training dataset consisting of 4 images per city (72 images in total). The original validation set has $500$ images, containing images from 3 different cities. We take $10$ images per city, resulting in a total of $30$ images. We convert the pixel-dense segmentation labels to patch-level classes by selecting the most common class within each patch. We evaluate the quality of representations by measuring pixel accuracy and segmentation mIoU. As the ViTs that we used take as inputs images of size $224\times 224$, we tile the images of size $1024\times 2048$ into $256\times 256$ tiles and treat each tile as a separate image. The tiles are further resized to $224\times 224$. We store patch representations for all images (training and validation) and ViTs.
Fair1M [29] is a satellite imagery dataset designed for fine-grained object detection. Note that none of the ViTs we tested were pretrained on satellite imagery (even DINOv2, as far as we can tell). The objects in FAIR1M are annotated in 5 supercategories: airplane, ship, vehicle, court, and road, and 37 fine-grained categories (types of airplanes, types of ships, etc.). The annotations are rotated bounding boxes (without pixel precision). To the best of our knowledge, all object instances appear on only one image. The images have varying sizes, usually larger than $1000\times 1000$px. We cropped them to $224\times 224$px tiles and kept only 8 images per each fine-grained category to ensure a minimum of 8 instances for each category. Tab. 3 in Appendix lists dataset statistics. We have 295 images with 196 patches each for ViT-B/16 models, and 256 patches each for DINOv2.
Tracking Dataset. For the tracking experiments, we take track-validation set of MOT 2020 images subset of BDD-100K dataset [32] and extract patch representations from each frame. We use ROIAlign [13] to pool features of each object instance based on its ground truth bounding box. For a given time difference $\Delta$ between two frames and for each frame at time step $t$, we retrieve the object with the closest embedding vector from the frame at time step $t+\Delta$.
4 k-NN Fails for MIMs: Proposed Solution
@@
In this section, we tackle a challenge with MIM-based methods in k-NN tasks. The problem revolves around features that have significantly higher variance, compared to other features. We suggest a solution for this issue and then, in Sec. 6, we take a closer look at what information these features hold.
Issue Discovery: k-NN vs. Linear Probing. To introduce the issue that we have noticed, we conduct the following experiment: We use two simple classifiers to analyze the underlying representations: k-NN with $k=1$ and fitting a linear softmax classifier. Both are trained on the patch representations of the few-shot training set. The motivation for employing these two basic methods is to understand whether the patches of a given object category cluster together or are linearly separable from other object categories in the representation space. For MAE, we follow recommendations from [15] (which on its turn refers to [10]) and apply batch normalization to the extracted features before the linear layer achieving 2.5×2.5$\times$2.5 × better results for linear probing with BatchNorm. We do not use BatchNorm for SimMIM, as adding it worsens the performance. The results are demonstrated in Fig. 1. We observe that MAE and SimMIM perform significantly worse for k-NN compared to linear probing. Since both methods employ pixel-level reconstruction loss, we hypothesized that something may be wrong with the obtained representations and proceeded to analyze their performance further.
@@
Throughout our experiments we discovered that MAE’s patch embeddings perform reasonably well with linear probing, but fail with k-NN. We hypothesize that the patch embeddings of MAE might have high variance in some dimensions which could drastically increase the distances among same-category patches and therefore harm k-NN, while not affecting the performance of linear models.
@@
(a) Variances of all (768) features of embeddings of various ViTs sorted in decreasing order. Both axis are logarithmic.
@@
Why MIM-based models have poor k-NN performance? We computed the variance for each of the features for all models. In Fig. 1(a), we plot the variances of all 768 features in a decreasing order. We observe that the variances throughout all features of Supervised ViT (and DINO) are relatively uniform. However, for MAE and SimMIM, there are several features with very high variances and there is a long tail of close to zero variance features.
Simple Remedy: Improving k-NN performance for MAE. We removed $m$ features with the highest variances and measured few-shot segmentation performance using k-NN and linear probing with the shortened embeddings. As seen in Fig. 1(b), with just $m=10$ features removed, MAE’s k-NN performance jumped from $0.058$ to $0.170$, without sacrificing the performance of linear probing. The performance of k-NN kept increasing up to $0.295$ at $m=200$. Afterwards, the scores for both k-NN and linear probing started to decline. This finding implies that around a quarter of features of MAE’s embeddings are not useful for patch-level image segmentation for neither linear models nor k-NNs. On the other hand, these features comprise nearly all of the variance of the embeddings.
Connections to Dimensional Collapse. The findings mentioned above align with that in [16], which asserts that some contrastive learning algorithms experience dimensional collapse. This occurs when the representation space fails to utilize its entirety, instead maps all images to a lower-dimensional subspace. Our findings indicate significantly higher variances in some features, suggesting that the representations lie in the space spanned by these high-variance features. We show that by removing these features, the performance of MAE for downstream tasks improves for k-NN and does not deteriorate for linear probing. These findings are somewhat orthogonal to that of [16]. This requires further analysis and is beyond the scope of this work.
5 Results
5.1 Experiments on CityScapes
The size of the training set matters for linear models, less for k-NN. In Fig. 1 we illustrate the performance of k-NN and linear classifiers for varying sizes of training dataset. We first observe that the linear models generally achieve better results than k-NN, especially with more training data. However, for the DINO versions and the supervised ViT, k-NN and linear classifiers demonstrate comparable performance. Instead, MAE representations yield bad k-NN performance. While its linear results are only slightly inferior to its DINO counterpart, the k-NN classifier leads to a remarkable about 4 times worse performance.
We applied three sets of corruptions to the same small subset of the validation set of Cityscapes. First, we blur out the images by applying four sizes of the blur kernel: $10\times 10$, $20\times 20$, $30\times 30$ and $40\times 40$. Second, we add four different levels of Gaussian noise (with mean 00 and standard deviations $10,20,30,$ and $40$). At last, we add frequency-based random noise (with $40$ standard deviation) to the images the following way. We generate a Gaussian noise of the same shape as the source image, convert it into the frequency domain, keep the frequencies only inside one of the four narrow bands, convert the noisy image back to image pixels, and add it to the original image. Fig. 10 in Appendix visualizes this process. We feed these degraded images to the ViTs and check the robustness of the methods for the different types and levels of degradations. The first row of Fig. 4 demonstrates the results of k-NN for blurred, additive Gaussian noise and added frequency-based random noise for each color, respectively.
One can claim that the robustness to various degradations may be attributed to the augmentations used during training. Models that employ color enhancement augmentations are assumed to exhibit higher resilience against those types of degradations (blur, Gaussian noise, etc). However, according to [15] color jittering-based augmentation degrades its performance. This suggests a potential trade-off between performance and robustness to degradations. Due to computational limitations, further investigation of this is left for future work.
Figure 4: Robustness of linear models for segmentation on Cityscapes. For every model (differing by colors) we have five dots, each representing one level of degradation. y-axis shows the mIoU on corrupted images. The top dots in vertical lines correspond to the original images. Grey diagonal lines indicate levels of equal relative drop in mIoU.